OpenStudy (anonymous):

help with this question: solve with elimination. 6x+y=-10 3x-y=3

OpenStudy (anonymous):

can you see that if you add these toghether the y variable will cancel out?

OpenStudy (anonymous):

yes i did that but after that is confusing

OpenStudy (anonymous):

Well, it leaves other stuff - try to add them together.

OpenStudy (anonymous):

well after adding i get, and hopefully you get: 9x=-7

OpenStudy (anonymous):

Which means that the x is something, and when you know x - voila!

OpenStudy (anonymous):

wait i eliminated the y's then i got: 6x=-12 and 3x=3

OpenStudy (anonymous):

how did you get that? if you add 6x+3x you get 9x right?

OpenStudy (anonymous):

you add the 6x and the 3x and the -12 and the 3?

OpenStudy (anonymous):

You didn't eliminate them, you added them together - do that with the rest as well:)

OpenStudy (anonymous):

where is the negative 12 coming from?

OpenStudy (anonymous):

6x+y=-10 3x-y=3 + ------------- 9x=-7

OpenStudy (anonymous):

oh sorry the -10 was originally a -12, i typed it out wrong

OpenStudy (anonymous):

okay anyway this is what you get: 6x+y=-12 3x-y=3 + ------------- 9x=-9

OpenStudy (anonymous):

so it would be: 9x=-9, right

OpenStudy (anonymous):

so x=-1

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

now, just plug that into any of the equations and solve for y

OpenStudy (anonymous):

i get 2

OpenStudy (anonymous):

but the answer key says it should be -1 and -6

OpenStudy (anonymous):

How do you get 2 then?

OpenStudy (anonymous):

x=-1 which is correct.

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

yes but i need a -6 too

OpenStudy (anonymous):

Just write the equations again. Then you write -1 where all the x-es are (its good to use () ). Then solve it as a regular problem.

OpenStudy (anonymous):

ok, thanks i got it

OpenStudy (anonymous):

You can use whichever problem you want, but one is usually easier to solve. Woohoo!:D

OpenStudy (anonymous):

Its also smart to check if both problems add up once you've got both of them.

OpenStudy (anonymous):

thanks